OpenStudy (anonymous):

Lucas bought some plumbing equipment for $37.38. The price he paid included a sales tax of 5%. What was the price of the plumbing equipment before the tax was added? A. $35.60 B. $36.33 C. $37.33 D. $39.25

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

dude i need help like bad >_<

OpenStudy (igreen):

Make an equation

OpenStudy (anonymous):

my mind hurts, it been a long day, so idk what to do

OpenStudy (igreen):

x + (0.05x) = 37.38

OpenStudy (igreen):

1 + 0.05 = ?

OpenStudy (anonymous):

1.05

OpenStudy (igreen):

Yep, so we have: 1.05x = 37.38 Now divide 1.05 to both sides

OpenStudy (igreen):

37.38 / 1.05 = ?

OpenStudy (anonymous):

35.6

OpenStudy (anonymous):

kk thx

OpenStudy (igreen):

Yep, that's our answer.